OpenStudy (destinykiara99):

Solve x^2 +12x + 6 = 0 using the completing-the-square method. a. x = negative six plus or minus the square root of thirty b. x = six plus or minus the square root of thirty c. x = negative six plus or minus the square root of six d. x = six plus or minus the square root of six

OpenStudy (welshfella):

divide the coefficient of x ( that is 12) by 2 to give (x + 6)^2 x^2 +12x + 6 = 0 (x + 6)^2 - 36 + 6 = 0 (x + 6)^2 - 30 = 0 (x + 6)^2 = 30 can you continue?